# Spectral Concentration at the Edge of Stability: Information Geometry of Kernel Associative Memory

### v3

@@ 1 Introduction
Although our focus is on associative memory, the geometric mechanisms revealed here, specifically Spectral Concentration and Dual Equilibrium, may offer broader insights into the generalization capabilities of over-parameterized neural networks, where similar phenomena of “feature learning” and “sharpness selection” have been reported.
@@ 3.1 Effective Dimensionality of Memory
Observation 1 (L-Shaped Spectrum): Numerical analysis reveals Fig. 1 reveals a striking contrast in the spectral structures. In the local regime (blue dashed line), the spectrum is flat ($\lambda_{k}\approx\lambda_{1}$), indicating a diffuse distribution of information where no dominant direction exists. In contrast, on the Ridge (red solid line), the spectrum exhibits a characteristic “L-shape”: the normalized eigenvalues drop by orders of magnitude immediately after the leading mode ($\lambda_{2}\ll\lambda_{1}$), yet the tail remains non-zero and flat ($\lambda_{k>1}>0$). This confirms that the Ridge achieves stability via Spectral Concentration, effectively compressing noise into a single dominant mode while preserving the dimensionality required for memory capacity. Quantitatively, while the local regime exhibits $d_{\text{eff}}(G)\approx P$ (full utilization of dimensions), the Ridge collapses this to $d_{\text{eff}}(G)\approx 1.5\sim 5.0$, depending on the load. This drastic reduction confirms that the network performs optimal compression, retaining only the essential degrees of freedom.
@@ 4.1 Maximal Curvature at the Ridge
The local curvature of the statistical manifold is quantified by the largest eigenvalue of the FIM, $\lambda_{\max}(G)$. In gradient descent dynamics, the stability of the update steps is constrained by this curvature; excessively sharp curvature can destabilize the learning process (Edge of Stability [8]).
@@ 5 Dual Geometry and Force Antagonism
Observation 3 (Dual Balance): Numerical experiments (Fig. 3) demonstrate a striking contrast on the Ridge:
@@ 5.2 Geometric Derivation of Antagonism
The decomposition $\nabla L=\bm{F}_{d}+\bm{F}_{i}$ can be understood as splitting the gradient into the “signal” component and the “curvature correction” component. The condition $\nabla L\to\bm{0}$ (equilibrium) in Euclidean space implies $\bm{F}_{d}\approx-\bm{F}_{i}$. However, the condition $\tilde{\nabla}L\to\bm{0}$ in Riemannian space is much stronger and more fundamental. The Ridge represents a state where the system has found a Dual Equilibrium: a point where the intrinsic error is minimized despite the extreme curvature of the ambient space.
@@ 6.1 Self-Organized Criticality in Memory
Our finding that the optimal memory state lies at the brink of information collapse (Fig. 2) aligns with the hypothesis of Self-Organized Criticality (SOC) [9] in biological neural networks. The brain is thought to operate at a critical point to maximize information transmission and dynamic range. Similarly, our KLR Hopfield network tunes itself to the Edge of Stability, maximizing the curvature of the statistical manifold to create deep attractors, while stopping just short of the singular region where memory capacity vanishes. This suggests that SOC may be a universal requirement for high-capacity associative memory.
@@ 7 Conclusion
This geometric framework unifies the apparently contradictory observations of instability (high curvature, Edge of Stability) and stability (deep attractors, Dual Balance). The Ridge of Optimization is not merely a parameter sweet spot, but a manifestation of Self-Organized Criticality in the information geometry of neural networks. The network tunes itself to the brink of instability, maximizing curvature to carve the deepest possible attractor basins, while maintaining dual equilibrium to ensure learning convergence.

### v4

@@ 1 Introduction
Although our focus is on associative memory, the geometric mechanisms revealed here, specifically Spectral Concentration and Dual Equilibrium, may offer broader insights into the generalization capabilities of over-parameterized neural networks, where similar phenomena of “feature learning” and “sharpness selection” have been reported. Rather than introducing an entirely new learning principle, this work reinterprets the Ridge of Optimization as an emergent boundary where existing geometric quantities, such as the Fisher Information Matrix and natural gradient dynamics, become maximally expressed.
@@ 3.1 Effective Dimensionality of Memory
Observation 1 (L-Shaped Spectrum): Numerical analysis (Fig. 1) reveals a striking contrast in the spectral structures. In the local regime (blue dashed line), the spectrum is flat ($\lambda_{k}\approx\lambda_{1}$), indicating a diffuse distribution of information where no dominant direction exists. In contrast, on the Ridge (red solid line), the spectrum exhibits a characteristic “L-shape”: the normalized eigenvalues drop by orders of magnitude immediately after the leading mode ($\lambda_{2}\ll\lambda_{1}$), yet the tail remains non-zero and flat ($\lambda_{k>1}>0$). This confirms that the Ridge achieves stability via Spectral Concentration, effectively compressing noise into a single dominant mode while preserving the dimensionality required for memory capacity. Quantitatively, while the local regime exhibits $d_{\text{eff}}(G)\approx P$ (full utilization of dimensions), the Ridge collapses this to $d_{\text{eff}}(G)\approx 1.5\sim 5.0$, depending on the load. This drastic reduction confirms that the network performs optimal compression, retaining only the essential degrees of freedom.
@@ 4.1 Maximal Curvature at the Ridge
The local curvature of the statistical manifold is quantified by the largest eigenvalue of the FIM, $\lambda_{\max}(G)$. In gradient descent dynamics, the stability of the update steps is constrained by this curvature; excessively sharp curvature can destabilize the learning process, a phenomenon known as the Edge of Stability [8]. We refer to the Ridge as an “Edge of Stability” in this context, not as a new dynamical phase, but as a geometric boundary characterized by the maximal sensitivity encoded in the Fisher Information Matrix.
@@ 5 Dual Geometry and Force Antagonism
Observation 3 (Dual Equilibrium): Numerical experiments (Fig. 3) demonstrate a striking contrast on the Ridge:
@@ 5.2 Geometric Derivation of Antagonism
The decomposition $\nabla L=\bm{F}_{d}+\bm{F}_{i}$ can be understood as splitting the gradient into the “signal” component and the “curvature correction” component. The condition $\nabla L\to\bm{0}$ (equilibrium) in Euclidean space implies $\bm{F}_{d}\approx-\bm{F}_{i}$. However, the condition $\tilde{\nabla}L\to\bm{0}$ in Riemannian space is much stronger and more fundamental. The Ridge represents a state where the system has found a Dual Equilibrium: a point where the intrinsic error is minimized despite the extreme curvature of the ambient space. This dual equilibrium does not introduce a new equilibrium concept, but highlights the coexistence of instability in the Euclidean parameter space and stability in the information-geometric (Riemannian) sense.
@@ 6.1 Self-Organized Criticality in Memory
Our finding that the optimal memory state lies at the brink of information collapse (Fig. 2) is reminiscent of the hypothesis of Self-Organized Criticality (SOC) [9] in biological neural networks, in the sense that the system naturally tunes itself to a critical boundary. The brain is thought to operate at a critical point to maximize information transmission and dynamic range. Similarly, our KLR Hopfield network tunes itself to the Edge of Stability, maximizing the curvature of the statistical manifold to create deep attractors, while stopping just short of the singular region where memory capacity vanishes. This suggests that SOC may be a universal requirement for high-capacity associative memory.
@@ 7 Conclusion
This geometric framework unifies the apparently contradictory observations of instability (high curvature, Edge of Stability) and stability (deep attractors, Dual Equilibrium). The Ridge of Optimization is not merely a parameter sweet spot, but a manifestation of Self-Organized Criticality in the information geometry of neural networks. The network tunes itself to the brink of instability, maximizing curvature to carve the deepest possible attractor basins, while maintaining dual equilibrium to ensure learning convergence.